# A Guide to Tracking Phylogenies in Parallel and Distributed Agent-based Evolution Models

### v1

@@ Methods
We used triplet distance to assess the overall quality of reconstruction (critchlow1996triples). This approach considers all possible three-leaf subsets of a phylogeny, and reports the fraction of triplets with topology mismatching the corresponding triplet in a reference tree. Triplet distance ranges from 0.0 (between identical trees) to 0.5 (between random trees) to a hypothetical maximum of 1.0 This measure requires two conditions to be satisfied: that (1) phylogenies are rooted [^2] Quartet distance, which considers topologies of four-leaf subsets, is required in the unrooted case (estabrook1985comparison). and (2) taxa correspond one-to-one between reconstruction and reference. Both conditions are satisfied; note that hereditary stratigraphy inherently produces rooted trees, a benefit of differentia time points being assigned relative to an explicit generation zero.
To discern error arising from unresolved (as opposed to incorrect) reconstruction, we included a second reconstruction quality measure: inner node loss. This metric quantifies the difference between the number of inner nodes present in the reference tree versus the reconstruction. It serves as a precision measure, designed to assess the amount of phylogenetic detail lost due to artifactual polytomization. Inner node loss ranges from 0.0 (for reconstruction with as many inner nodes as reference) to a maximum of 1.0 (reconstruction is a pathological star phylogeny with only one inner node). Note that a negative inner node loss might be measured if the reconstruction contains more inner nodes than the reference, owing to erroneous overresolution. This might occur, for instance, due to the inherent inability of bit-width differentia to represent node degrees higher than bifurcation. For bit-differentia configurations, as remarked above, this inner node loss measure assumes a very specific interpretation: artifactual polytomies occur exclusively when annotation records share all differentia in common. In this case, identical annotations are polytomized as leaves of a single inner node.
To further discern error from unresolved versus incorrect reconstruction, we on occasion distinguish an alternate “lax” triplet distance from the “strict” triplet distance measure described above. Lax triplet distance differs from strict triplet distance in that it does not penalize triplets that mismatch on account of polytomy. This measure is useful in isolating incorrect reconstruction from unresolved reconstruction. However, care should be taken in interpreting lax triplet distance, as the pathological “star” tree case where all leaves descend directly from the root in one large polytomy would measure zero reconstruction error under the lax triplet distance measure. As such, where used strict triplet distance is also reported. Where not specified, triplet distance refers to the strict measure.
@@
Figure 3 shows example reference phylogenies, corresponding reconstructions, and quality metrics in practice. The top panel shows trials from the drift treatment, which exhibits high phylogenetic richness, and the bottom panel shows trials from the plain treatment, which exhibits low phylogenetic richness. (Note that, for legibility, time axes of phylogenies are log-scaled, which somewhat reduces the apparent visual distinction of high-versus-low phylogenetic richness.) Each panel compares reconstruction results under steady (left) and recency-proportional (right) instrumentation. Particularly high incidence of node loss (green triangle) can be seen under steady retention. Correspondingly, many large polytomies can be seen in the example steady-retention reconstructions (blue overlaid dendrogram) compared to the corresponding reference phylogeny (orange underlaid dendrogram). Under the plain treatment, steady retention leads to particularly high (almost complete) inner node loss, and — correspondingly — triplet distance similarity (purple dots) is very poor. Despite high inner node loss under the drift treatment, triplet distance remains low under steady retention. Across example cases shown, tilted retention enjoys triplet distance and inner node loss comparable to or better than steady retention. The main discussion will return to explore this question of steady-versus-tilted retention in greater rigor and depth.
@@
Comparisons of reconstruction quality considered both statistical significance (the likelihood an observed difference between treatments might have occurred by chance) and effect size (the magnitude of distinction between treatments relative to outcome variabilities). We used nonparametric methods for both analyses. By assessing effect sign, size, and significance across treatment conditions, we were able to describe the extent and consistency with which one instrumentation approach outperformed others.
@@
We pair effect-size analysis with Mann-Whitney U testing in order to assess evidence that significant differences exist between reconstruction quality under different conditions (mann1947on). As our goal was to screen for possible effects, rather than establish the veracity of any one effect, we did not correct for multiple comparisons in assessing statistical significance.
When determining the best- and worst-performing among three or more hereditary stratigraphy approaches, we use a nonparametric skimming procedure provided by the pecking Python library (moreno2024pecking). The procedure first applies a Kruskal-Wallis H-test to determine if there is evidence for significant variation among the sample groups. If this test fails, the groups are no best- or worst-performing group or groups are identified. If a significant difference is found, observation ranks are calculated and sample groups are sorted in order of mean rank. To discern the lowest-ranked group(s), successive Mann-Whitney U-tests are performed between the lowest-rank group and successively higher-ranked groups, adjusting the significance level (‘alpha‘) for multiple comparisons according to a sequential Holm-Bonferroni program. The overall lowest-ranked group and subsequent groups tested before encountering the first significantly differing group are taken as the best-performing (given that the metric in question is an error measure). This approach, therefore, identifies the set of lowest-ranked groups that are statistically indistinguishable amongst themselves. A similar procedure is used to identify a set of highest-ranked groups.
@@
In this section, we report annotate-and-reconstruct experiments comparing phylogeny reconstruction quality obtained across possible hereditary stratigraphy approaches. These experiments seek to establish a holistic, evidence-driven synopsis of each approach’s suitability across experimental use cases. The following section, “A Practicioner’s Guide to Hereditary Stratigraphy,” then synthesizes findings to suggest guidelines for selecting appropriate methods to apply in practice.
This section delves into three primary aspects of hereditary stratigraphy methodology:
@@
Surface vs. Column Implementaiton
@@
Recall that surface- and column-based annotation implementations differ in how differentia are organized within a hereditary stratigraphy annotation. Surface-based implementation takes a lower-level approach that simplifies accessioning of new differentia and ensures full use of available memory space but sacrifices some control over the precise temporal distribution of retained differentiae. Both implementations support tilted and steady retention policies.
@@
Under tilted retention, triplet distance (a measure of reconstruction accuracy) is equivalent or improved (in 14 / 48 scenarios) with surface-based implementation. Inner node loss (a measure of reconstruction precision) improves in some scenarios and worsens in others. Notably, shown in Supplementary Figure 9, inner node loss improvement is seen in all treatments with byte-width differentiae, which are mechanistically more prone to create artifactual unresolved polytomies that drive inner node loss. In sum, reconstruction quality of surface-based tilted retention can be considered equivalent or superior across the board to that of column-based implementation.
@@
In contrast, under steady retention, surface-based implementation achieves worse triplet distance reconstruction in 11 / 48 scenarios and better triplet distance in no scenario. Similarly, inner node loss is worse in 23 / 48 scenarios, including 4 / 12 byte-differentiae scenarios, and better in no scenario. So, reconstruction quality of surface-based steady retention is equivalent or inferior to column-based implementation.
@@
Why does the surface-based approach benefit reconstruction under one retention policy but not the other? Compared to column-implementation, surface-implementation allows the tilted algorithm to retain more differentia within available annotation space. Whereas the write-only design of the surface-based approach guarantees full use of buffer space, the column-based tilted retention holds space in reserve on account of discretization effects in tuning retention density. In contrast, column-based steady retention makes full use of available space, giving the surface-based approach no advantage in this regard. Although both approaches prune differentiae through comparable strided decimation procedures, the column implementation more systematically drops decimated differentiae from back to front. This process ends up preserving more recent differentiae which, as we will see in the next set of experiments comparing steady and tilted retention, tends to benefit reconstruction quality.
@@
Although surface-based implementation involves a trade-off between performance reconstruction quality in the steady case, it is notable that surface-based implementation provides uncompromised improvement in both runtime performance and data quality.
@@
Steady and tilted retention differ in the composition of differentia checkpoint records maintained within hereditary stratigraphy annotations. Recall that steady policy spaces retained differentia evenly across history, while tilted policy retains more of more recent differentia. The question of which policy gives higher quality reconstruction boils down to where precision in discerning the timing of branching events is most useful to resolving evolutionary history. In the interest of even footing, we report results for each retention policy using its best-performing implementation, as established above: steady policy uses column implementation and tilted policy uses surface implementation. We also consider a hybrid policy, which splits surface buffer space evenly between steady and tilted retention.
@@
Figure 5 overviews how reconstruction quality differs by retention policy across use case scenarios. Across the board, steady policy yields phylogenetic reconstruction with heaviest inner node loss. In contrast, tilted policy exhibits among the lowest inner node loss in all but six surveyed scenarios — including all byte-differentia trials (Supplementary Figure 10). The hybrid policy has lower inner node loss in those six scenarios, which all involve evolutionary conditions with high phylogenetic richness.
@@
Steady policy also consistently produces the worst triplet distance error in lower-phylogenetic-richness plain and mild structure evolutionary scenarios. However, in scenarios with high phylogenetic richness, triplet distance error under steady retention fares better. With rich ecological/spatial structure, triplet distance reconstruction error is largely indistinguishable among retention policies. Further, steady policy triplet error actually significantly outperforms tilted policy in several instances under drift conditions. In nearly all of these instances, though, hybrid policy triplet error performs comparably to steady retention. In low phylogenetic richness plain and mild-structure scenarios, hybrid triplet error is comparable to tilted error in 9 / 24 scenarios and outperformed by tilted policy in 15 / 24 scenarios.
@@
Across the inner node loss and triplet error quality measures, tilted retention frequently performs best and steady retention frequently performs worst. However, tilted retention has worse triplet error in some scenarios with high phylogenetic richness. Hybrid retention performs more consistently across evolutionary scenarios. It exhibits consistently intermediate levels of inner node loss that, in absolute terms, tend not to be far off from tilted retention. Triplet error for hybrid retention is often comparable to the better-performing of steady and tilted retention, or at least intermediate between them. For greater detail of steady vs. tilted reconstruction quality outcomes broken down by treatment condition, see Supplementary Figure 9.
@@
Figure 6: How does retention policy affect the structure of inner node loss? Reconstruction node count densities for 256-bit size, byte differentia treatments. Histograms depict internal node count relative to reference trees (dashed line), binned by time ago (ranging from most recent to most ancient). Facet columns differentiate evolutionary regimes and facet rows correspond to steady, hybrid, and tilted retention policies. Owing to its non-prioritization to retain recent differentiae, steady policy loses nearly all resolution over recent (c. 100 generations) evolutionary history. Tilted policy, on the other hand, reconstructs the highest proportion of recent history. Hybrid policy does not suffer the catastrophic inner node loss over recent events seen under steady policy, but also reconstructs somewhat fewer recent inner nodes compared to pure tilted policy.
@@
Differences in retention of recent differentia explain the substantial advantage of tilted policy in use cases with low phylogenetic richness. In such scenarios, frequent selective sweeps concentrate phylogenetic history over very recent history, meaning that lineage-branching events giving rise to a contemporary extant population occurred over a relatively short period of time. Discerning these events, therefore, requires densely packed differentia checkpoints over recent history. Otherwise, in the worst case, taxa would jump from all sharing the same lineage marker at one checkpoint to all having distinct checkpoint markers — resulting in a catastrophic unresolved polytomy. Steady retention maintains a uniform gap size between retained differentiae that grows linearly with generations elapsed, meaning that very recent history is sparsely covered, if at all. The consequences of this deficiency can be seen in Figure 6, which compares the density of reconstructed nodes to ground truth. Reconstructions from steady policy are entirely missing branching events over the prefatory hundred or so generations. Figures 7(a) and 7(b) show reconstruction outcomes resulting under this inner node loss. Under steady policy, very high levels of unresolved reconstruction occur in scenarios with low phylogenetic richness, particularly for recent branching events. Example reconstructions exhibiting catastrophic comb polytomies characteristic of steady reconstruction of low-richness phylogenies can be seen in 3.
@@
Figure 7: How does differentia width affect reconstruction quality? Top panels show reconstruction outcomes for phylogenetic branching events (Figure 2) under bit- and byte-width differentiae, respectively. Outcomes are binned by time ago (ranging from most recent to most ancient). Bottom panel compares reconstruction quality metrics between reconstructions from annotations with bit- and byte-width differentiae. Color coding indicates significance, with red indicating better bit-width differentia performance and green indicating better byte-width differentia performance Byte-width differentia annotation consistently underperforms bit-width differentia in inner node loss and strict triplet distance measures. However, byte-width differentia outperform bit-width differentia in the lax triplet distance measure, which does not penalize polytomy triplets, i.e., isolating incorrect reconstruction from unresolved reconstruction (Figure 2). See Supplementary Figure 11 for listing of effects by sensitivity analysis condition.
@@
We now turn to consider the role of differentia size in hereditary stratigraphy reconstruction quality. Intuitively, tuning differentia size would seem to trade-off between accuracy and precision. Larger differentiae reduce the probability of spurious collision, which falsely makes lineages appear more closely related than they actually are. Note, in particular, that reconstructions from bit-sized differentia estimate all history as bifurcating, because each checkpoint can only discern two distinct lineages. On the other hand, for fixed annotation size, widening differentia necessarily reduces differentia count. Thus, differentia size diminishes the granularity at which branching events can be dated.
@@
To ascertain the actual effects of differentia width on reconstruction quality, we performed annotate-and-reconstruct experiments across a variety of use case scenarios. These experiments used 256-bit-sized annotations, comprised of either 256-bit-sized differentiae or 32-byte-sized differentiae. Figure 7(c) overviews the relative performance of bit- and byte-width differentiae on triplet distance and inner node loss quality measures. (Supplementary Figure 11 presents these results in greater detail, showing reconstruction outcomes for each treatment condition surveyed.) However, byte-width differentia consistently produces reconstructions with lower inaccuracy — as measured by lax triplet distance, which does not penalize unresolved reconstruction. Byte-width’s lower incidence of incorrect reconstruction outcomes is apparent in Figures 7(a) and 7(b), which assess rates of correct, incorrect, and unresolved reconstruction outcomes across evolutionary history. Unlike bit-width annotations, byte-width methods produce almost no incorrect reconstruction outcomes. However, owing to unresolved byte-width outcomes, bit-width annotations nonetheless have a higher rate of correct outcomes. Figure 7(c) indeed confirms that, in nearly all cases, bit-width differentiae produce more informative depictions of phylogenetic history, as measured by strict triplet distance.
@@
Figure 8: How does scale impact reconstruction quality? Each dot is a Cliff’s Delta effect size of downscale-vs-upscale comparison under one sensitivity analysis condition. Color coding indicates a significant outcome (Mann-Whitney U). In order, rows show scaling outcomes for increasing population size, population subsample size (i.e., tree tip count), and both these factors. For tilted and hybrid policies, scaling had minimal effects on triplet distance and no systematic effect on inner node loss. Inner node loss worsened when scaling sample size, with and without scaling population size. Scaling effects were more variable under steady retention policy. See Supplementary Figures 13, 12 and 14 for listing of effects by sensitivity analysis condition.
@@
Inner node loss, under hybrid and tilted retention, is generally also robust to population scaling. For these policies, inner node loss worsens only in use cases with byte-width differentiae (Supplementary Figures 13 and 14). In other cases, inner node loss actually often improves with population scale. Sample size scaling, however, tends to aggravate inner node loss — whether in isolation or in concert with population size scaling. As an exception, we do not see inner node loss worsen with larger sample size in some cases with byte-width differentia under tilted and hybrid policies (Supplementary Figures 13 and 14). Likewise, inner node loss remains stable for tilted retention under drift conditions (Supplementary Figure 14).
A Practicioner’s Guide to Hereditary Stratigraphy

### v2

@@ Methods
We used triplet distance to assess the overall quality of reconstruction (critchlow1996triples). This approach considers all possible three-leaf subsets of a phylogeny, and reports the fraction of triplets with topology mismatching the corresponding triplet in a reference tree. Triplet distance ranges from 0.0 (between identical trees) to 0.5 (between random trees) to a hypothetical maximum of 1.0. The triplet distance measure requires that phylogenies are rooted (estabrook1985comparison). Conveniently, hereditary stratigraphy produces rooted trees, owing to differentia time points being assigned relative to an explicit generation zero.
To discern error arising from unresolved (as opposed to incorrect) reconstruction, we included a second reconstruction quality measure: inner node loss. This metric quantifies the difference between the number of inner nodes present in the reference tree versus the reconstruction. It serves as a precision measure, designed to assess the amount of phylogenetic detail lost due to artifactual polytomies. Inner node loss ranges from 0.0 (for reconstruction with as many inner nodes as reference) to a maximum of 1.0 (reconstruction is a pathological star phylogeny with only one inner node). Note that a negative inner node loss might be measured if the reconstruction contains more inner nodes than the reference, owing to erroneous overresolution. This might occur, for instance, due to the inherent inability of bit-width differentia to represent node degrees higher than bifurcation. For bit-differentia configurations, as remarked above, this inner node loss measure assumes a very specific interpretation: artifactual polytomies occur exclusively when annotation records share all differentiae in common. In this case, identical annotations are polytomized as leaves of a single inner node.
To further discern error from unresolved versus incorrect reconstruction, we occasionally distinguish an alternate “lax” triplet distance from the “strict” triplet distance measure described above. Lax triplet distance differs from strict triplet distance in that it does not penalize triplets that mismatch on account of polytomy. This measure is useful in isolating incorrect reconstruction from unresolved reconstruction. However, care should be taken in interpreting lax triplet distance, as the pathological “star” tree case where all leaves descend directly from the root in one large polytomy would measure zero reconstruction error under the lax triplet distance measure. As such, where used, strict triplet distance is also reported. Where not specified, triplet distance refers to the strict measure.
@@
Figure 3 shows example reference phylogenies, corresponding reconstructions, and quality metrics in practice. The top panel shows trials from the drift treatment, which exhibits high phylogenetic richness, and the bottom panel shows trials from the plain treatment, which exhibits low phylogenetic richness. (Note that, for legibility, time axes of phylogenies are log-scaled, which somewhat reduces the apparent visual distinction of high-versus-low phylogenetic richness.) Each panel compares reconstruction results under steady (left) and recency-proportional (right) instrumentation. Particularly high node loss (green triangle) can be seen under steady retention. Correspondingly, many large polytomies can be seen in the example steady-retention reconstructions (blue overlaid dendrogram) compared to the corresponding reference phylogeny (orange underlaid dendrogram). Under the plain treatment, steady retention leads to very high (almost complete) inner node loss, and — correspondingly — triplet distance similarity (purple dots) is very poor. Despite high inner node loss under the drift treatment, triplet distance remains low under steady retention. Across example cases shown, tilted retention enjoys triplet distance and inner node loss comparable to or better than steady retention. The main discussion will return to explore this question of steady-versus-tilted retention in greater rigor and depth.
@@
Comparisons of reconstruction quality considered both statistical significance (the likelihood an observed difference between treatments might have occurred by chance) and effect size (the magnitude of distinction between treatments relative to outcome variabilities). We used nonparametric methods for both analyses. By assessing effect sign, size, and significance across treatment conditions, we described the extent and consistency with which one instrumentation approach outperformed others.
@@
We pair effect-size analysis with Mann-Whitney U testing in order to assess evidence that significant differences exist between reconstruction quality under different conditions (mann1947on). As our goal was to screen for possible effects, rather than establish the veracity of any one effect, we did not correct for multiple comparisons in assessing statistical significance. Because incorrectly identifying a lack of difference between conditions would be more harmful than incorrectly identifying the presence of a difference, this approach is conservative in maintaining the sensitivity of our assays. Where we do detect a significant effect, we additionally report results after applying Bonferroni correction.
When determining the best- and worst-performing among three or more hereditary stratigraphy approaches, we use a nonparametric skimming procedure provided by the pecking Python library (moreno2024pecking). The procedure first applies a Kruskal-Wallis H-test to determine if there is evidence for significant variation among the sample groups. If this test fails, then no best- or worst-performing group or groups are identified. If a significant difference is found, observation ranks are calculated and sample groups are sorted in order of mean rank. To discern the lowest-ranked group(s), successive Mann-Whitney U-tests are performed between the lowest-rank group and successively higher-ranked groups, adjusting the significance level $\alpha$ for multiple comparisons according to a sequential Holm-Bonferroni procedure. The overall lowest-ranked group and subsequent groups tested before encountering the first significantly differing group are taken as the best-performing (given that the metric in question is an error measure). This approach, therefore, identifies the set of lowest-ranked groups that are statistically indistinguishable amongst themselves. A similar procedure is used to identify a set of highest-ranked groups.
@@
This section reports annotate-and-reconstruct experiments comparing phylogeny reconstruction quality obtained across possible hereditary stratigraphy approaches. These experiments seek to establish a holistic, evidence-driven synopsis of each approach’s suitability across experimental use cases. The following section, “A Practicioner’s Guide to Hereditary Stratigraphy,” then synthesizes findings to suggest guidelines for selecting appropriate methods to apply in practice.
We delve into three primary aspects of hereditary stratigraphy methodology:
@@
Surface vs. Column Implementation
@@
Recall that surface- and column-based annotation implementations differ in how differentiae are organized within a hereditary stratigraphy annotation. Surface-based implementation takes a lower-level approach that streamlines generation-to-generation updates and ensures full use of available memory space, but sacrifices some control over the temporal distribution of retained differentiae. Both implementations support tilted and steady retention policies.
@@
Surface vs. Column for Tilted Retention
Under tilted retention, triplet distance (a measure of reconstruction accuracy) significantly improves in $14/48$ scenarios under surface-based implementation, before correction for multiple comparisons. Column-based implementation significantly improves reconstruction accuracy in no scenarios when applying the Mann-Whitney U test. Looking at effect sizes via Cliff’s delta, surface-based approaches improve reconstruction quality with at least a small effect size in $33/48$ scenarios. Column-based approaches improve reconstruction quality with a small effect size in 1 /48 scenarios. Although no Mann-Whitney U tests from any individual scenario are significant after Bonferroni correction, a strong trend exists across scenarios of better reconstruction quality from surface-based implementation for tilted retention (exact Binomial test; $p<0.0001$).
@@
Inner node loss (a measure of reconstruction precision) significantly improves with surface-based implementation in some scenarios ($24/48$), and significantly worsens in others ($13/48$). All results remain significant after correction for multiple comparisons. Notably, shown in Supplementary Figure 9, inner node loss improvement under surface-based implementation is seen in all treatments with byte-width differentiae, which are more prone to create artifactual unresolved polytomies that drive inner node loss. In sum for tilted retention, reconstruction quality of surface-based annotations can be considered equivalent or superior across the board to column-based implementation.
Surface vs. Column for Steady Retention
In contrast, under steady retention, surface-based implementation achieves worse triplet distance in $11/48$ scenarios and better triplet distance in no scenarios. Only one Mann-Whitney U test remains significant after Bonferroni correction. However, a significant trend again exists across scenarios, with $25/48$ scenarios where surface-based implementation degrades triplet distance by at least a small effect size under Cliff’s delta (exact Binomial test; $p<0.0001$).
@@
Additionally, for steady retention, inner node loss is significantly worse under surface-based in $23/48$ scenarios and better in no scenarios. So, across the board, reconstruction quality of surface-based annotations under steady retention is equivalent or inferior to that of column-based implementation.
Surface vs. Column — Synthesis
@@
Why does the surface-based approach benefit reconstruction under one retention policy but not the other? Compared to column-based implementation, surface-based implementation allows the tilted algorithm to retain more differentiae within available annotation space. Whereas the write-only design of the surface-based approach guarantees full use of buffer space, column-based tilted retention wastes some space held in reserve due to algorithmic limitations. In contrast, column-based steady retention makes full use of available space, giving the surface-based approach no advantage in this regard. Although both the surface and column approaches prune differentiae through comparable strided decimation procedures, the column implementation more systematically drops decimated differentiae from back to front. This process ends up preserving more recent differentiae, which — as we will see in the next set of experiments — tends to benefit reconstruction quality.
@@
Although surface-based implementation involves a trade-off between performance and reconstruction quality in the case of steady retention, it is notable that surface-based implementation improves both runtime performance and data quality under tilted retention. Given this result, we next assess the extent to which tilted or steady retention would be preferred in practice across possible evolutionary scenarios.
@@
Recall that steady and tilted retention differ in the temporal composition of retained differentiae; steady policy spaces retained differentia evenly across history, while tilted policy biases toward retaining more recent differentiae. The question of which policy enables higher quality reconstruction boils down to where precision in discerning the timing of lineage branching events is most useful in resolving evolutionary history. To ensure even footing, we report results for each retention policy using its best-performing implementation, as established above. Steady policy uses column implementation and tilted policy uses surface-based implementation. We also consider a surface-based hybrid policy, which splits buffer space evenly between steady and tilted retention.
@@
Figure 5 overviews reconstruction quality by retention policy across use case scenarios. Across the board, steady policy yields phylogenetic reconstruction with heaviest inner node loss. In contrast, tilted policy exhibits among the lowest inner node loss in all but six surveyed scenarios — including all byte-differentia trials (Supplementary Figure 10). The hybrid policy has lower inner node loss in those six scenarios, which all involve evolutionary conditions with high phylogenetic richness.
@@
Steady policy also consistently produces the highest triplet distance error in lower-phylogenetic-richness plain and mild-structure evolutionary scenarios. However, in scenarios with high phylogenetic richness, triplet distance error under steady retention fares better. With rich ecological/spatial structure, triplet distance reconstruction error is largely indistinguishable among retention policies. Further, steady policy triplet error significantly outperforms tilted policy in several instances under drift conditions. In nearly all of these instances, though, hybrid policy triplet error performs comparably to steady retention. In low phylogenetic richness plain and mild-structure scenarios, hybrid triplet error is comparable to tilted error in $9/24$ scenarios and outperformed by tilted policy in $15/24$ scenarios.
@@
Across the inner node loss and triplet error quality measures, tilted retention frequently performs best and steady retention frequently performs worst. However, tilted retention has worse triplet error in some scenarios with high phylogenetic richness. Hybrid retention performs more consistently across evolutionary scenarios. It exhibits consistently intermediate levels of inner node loss that, in absolute terms, tend to be comparable to tilted retention. Triplet error for hybrid retention is often comparable to the better-performing of steady and tilted retention, or at least intermediate between them. For greater detail of steady vs. tilted reconstruction quality outcomes broken down by treatment condition, see Supplementary Figure 9.
@@
Figure 6: How does retention policy affect the structure of inner node loss? Reconstruction node count densities for 256-bit size, byte differentia treatments. Histograms depict internal node count relative to reference trees (dashed line), binned by time ago (ranging from most recent to most ancient). Facet columns differentiate evolutionary regimes and facet rows correspond to steady, hybrid, and tilted retention policies. Owing to its non-prioritization of retaining recent differentiae, steady policy loses nearly all resolution over recent (c. 100 generations) evolutionary history. Tilted policy, on the other hand, reconstructs the highest proportion of recent history. Hybrid policy does not suffer the catastrophic inner node loss over recent events seen under steady policy, but also reconstructs somewhat fewer recent inner nodes compared to pure tilted policy.
@@
Differences in retention of recent differentia explain the substantial advantage of tilted policy in use cases with low phylogenetic richness. In such scenarios, frequent selective sweeps concentrate phylogenetic history over very recent history, meaning that lineage-branching events giving rise to a contemporary extant population occur over a relatively short period of time. Discerning these events, therefore, requires densely packed differentia checkpoints over recent history. Otherwise, in the worst case, taxa would jump from all sharing the same lineage marker at one checkpoint to all having distinct checkpoint markers — resulting in a catastrophic unresolved polytomy. Steady retention maintains a uniform gap size between retained differentiae that grows linearly with generations elapsed, meaning that very recent history is thinly covered, if at all. The consequences of this deficiency can be seen in Figure 6, which compares the density of reconstructed nodes to ground truth. Reconstructions from steady policy are entirely missing branching events over the most recent generations. Figures 7(a) and 7(b) show reconstruction outcomes resulting under this inner node loss. Under steady policy, very high levels of unresolved reconstruction occur in scenarios with low phylogenetic richness, particularly for recent branching events. Example reconstructions exhibiting catastrophic comb polytomies characteristic of steady reconstruction of low-richness phylogenies can be seen in 3.
@@
Figure 7: How does differentia width affect reconstruction quality? Top panels show reconstruction outcomes for phylogenetic branching events (Figure 2) under bit- and byte-width differentiae, respectively. Outcomes are binned by time ago (ranging from most recent to most ancient). Bottom panel compares reconstruction quality metrics between reconstructions from annotations with bit- and byte-width differentiae. Color coding indicates significance, with red indicating better bit-width differentia performance and green indicating better byte-width differentia performance. Byte-width differentiae consistently underperform bit-width differentia in inner node loss and strict triplet distance measures. However, byte-width differentiae outperform bit-width differentia in the lax triplet distance measure, which does not penalize polytomy triplets, i.e., isolating incorrect reconstruction from unresolved reconstruction (Figure 2). See Supplementary Figure 11 for listing of effects by sensitivity analysis condition.
@@
We now consider the role of differentia size in hereditary stratigraphy reconstruction quality. Intuitively, tuning differentia size would seem to trade-off between accuracy and precision. Larger differentiae reduce the probability of spurious collision, which falsely makes lineages appear more closely related than they actually are. Note, in particular, that reconstructions from bit-sized differentia estimate all history as bifurcating, because each checkpoint can only discern two distinct lineages. On the other hand, for fixed annotation size, widening differentia necessarily reduces differentia count. Thus, differentia size diminishes the granularity at which branching events can be dated.
@@
To test the effects of differentia width on reconstruction quality, we performed annotate-and-reconstruct experiments across a variety of use case scenarios. These experiments used 256-bit-sized annotations, comprised of either 256-bit-sized differentiae or 32-byte-sized differentiae. Figure 7(c) overviews the relative performance of bit- and byte-width differentiae on triplet distance and inner node loss quality measures. (Supplementary Figure 11 presents these results in greater detail, showing reconstruction outcomes for each treatment condition surveyed.) However, byte-width differentia consistently produces reconstructions with lower inaccuracy — as measured by lax triplet distance, which does not penalize unresolved reconstruction. Byte-width’s lower incidence of incorrect reconstruction outcomes is apparent in Figures 7(a) and 7(b), which assess rates of correct, incorrect, and unresolved reconstruction outcomes across evolutionary history. Unlike bit-width annotations, byte-width methods produce almost no incorrect reconstruction outcomes. However, owing to unresolved byte-width outcomes, bit-width annotations nonetheless have a higher rate of correct outcomes. Figure 7(c) indeed confirms that, in nearly all cases, bit-width differentiae produce more informative depictions of phylogenetic history, as measured by strict triplet distance.
@@
Figure 8: How does scale impact reconstruction quality? Each dot shows a Cliff’s Delta effect size of downscale-vs-upscale comparison under one sensitivity analysis condition. Color coding indicates a significant outcome (Mann-Whitney U). In order, rows show scaling outcomes for increasing population size, population subsample size (i.e., tree tip count), and both these factors. For tilted and hybrid policies, scaling had minimal effects on triplet distance and no systematic effect on inner node loss. Inner node loss worsened when scaling sample size, with and without scaling population size. Scaling effects were more variable under steady retention policy. See Supplementary Figures 13, 12 and 14 for listing of effects by sensitivity analysis condition.
@@
Inner node loss, under hybrid and tilted retention, is generally also robust to population scaling. For these policies, inner node loss worsens only in use cases with byte-width differentiae (Supplementary Figures 13 and 14). In other cases, inner node loss improves with population scale. Sample size scaling, however, tends to aggravate inner node loss — whether in isolation or in concert with population size scaling. As an exception, we do not see inner node loss worsen with larger sample size in some cases with byte-width differentia under tilted and hybrid policies (Supplementary Figures 13 and 14). Likewise, inner node loss remains stable for tilted retention under drift conditions (Supplementary Figure 14).
A Practitioner’s Guide to Hereditary Stratigraphy